OpenStudy (anonymous):

7(-3x + 7) = -(-3x + -73)

OpenStudy (anonymous):

distribute the 7 on lhs and -1 on rhs and should be simpler.

OpenStudy (anonymous):

7(-3x+7)=-(-3x+-73) -21x+49=3x+73 -21x-3x=73-49 -24x=24 x=-1